OpenStudy (anonymous):

I really need help! 15x+3/27X-18*3X-2/5X+1 how would i do that?

OpenStudy (anonymous):

answer choices are: A. -2x-1/12x-15 B. 3 C. 1/3 12x-15/-2x-1

OpenStudy (zehanz):

Maybe you could put some brackets here and there to clarify the order of calculations?

OpenStudy (anonymous):

I tried everything I thought I could, My answer was D and it was wrong.

OpenStudy (zehanz):

I mean, it is not clear to me what the formula is. Is it:\[15x+\frac{ 3 }{ 27x }-18 \cdot 3x-\frac{ 2 }{ 5x }+1\]or is it:\[15x+\frac{ 3 }{ 27 }x=18 \cdot 3x -\frac{ 2 }{ 5 }x+1\] That makes a lot of difference!

OpenStudy (zehanz):

TYPO: second formula: replace "=" with minus sign...

OpenStudy (anonymous):

No sorry, hold on I'll draw it

OpenStudy (anonymous):